Jackson Autos has one employee. As of March 30, their employee had already earned $6,300. For the pay period ending April 15, their employee earned an additional $2,000 of gross wages. Only the first $7,000 of annual earnings are subject to FUTA of 0.6% and SUTA of 5.4%. The journal entry to record the employer's unemployment payroll taxes for the period ending April 15, would be:
Answer #1

Unemployment taxes would be charged up to the extent of $7,000 earnings.

Since $6,300 is already earned till March 30, the additional earnings of (7,000 – 6,300 =) $700 should be considered for taxes for the April 15 pay.

FUTA = $700 × 0.6% = $4.2

SUTA = $700 × 5.4% = $37.8

Total = 4.2 + 37.8 = $42

Journal

Date

Account titles and explanation

Ref.

Debit

Credit

April 15

Payroll taxes expense

$42.00

       FUTA Payable

$4.20

       SUTA Payable

$37.80

To record unemployment taxes to Federal government and State government.